OpenStudy (anonymous):

how do you do 16a^2+40q+25

OpenStudy (eyust707):

What is it you would like to do?

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

\[16a^2+40a+25\] to make it clearer.

OpenStudy (eyust707):

\[16a^2+40a+25 = (4a +5)(4a+5)= (4a+5)^2\]

OpenStudy (anonymous):

did you do it with like the trial and error method?

OpenStudy (anonymous):

@Kimberlina the method for this is calculate 25*16 = 400, now factorize the middle term so that sum of factors =40 and product of factors =400,i.e factors would be 20 and 20 , so you can write equations as \[16 a ^2 +20 a +20 a +25 = 4a (4a +5) + 5 (4a+5)\] =(4a+5)(4a+5) ...this method is also called middle term spilitting

OpenStudy (anonymous):

ohhhhhh okay i see ! @akash_809 .. This was actually very helpful .

OpenStudy (anonymous):

@Kimberlina thanks no problems